# WOMD-Reasoning: A Large-Scale Language Dataset for Interaction and Driving Intentions Reasoning

### v1

@@ 3.2 GPT-based Interaction Analysis
Categories Map Ego Agent Other Agents Interactions Intentions Total Scenes
Training Set 188k 268k 1,635k 287k 52k 2,430k 52k
Validation Set 41k 58k 341k 59k 11k 510k 11k
@@
4 Dataset
In this section, we provide detailed information about WOMD-Reasoning. We first present the organization and statistics of the dataset, then provide quality analysis based on case studies and preliminary human evaluations. Finally, we compare our dataset to previous ones to show the strengths of WOMD-Reasoning in diverse interactions, huge size and wide potential applications.
4.1 Dataset Details
Table 2 provides the statistics of Q&As in WOMD-Reasoning dataset, as well as in every Q&A class. Our Q&As can be classified into five categories: Map Environment, which includes the existence and category of intersections, existence and places of stop signs and crosswalks, and count of lanes, etc. Ego Agent’s Motion Status and Other (Surrounding) Agents’ Motion status, which describe each agent’s speed, acceleration, direction, related traffic light and sign, and relative position to the intersection center as well as to the ego agent. Interaction, which includes a summary of interactions between each surrounding agent and the ego agent, and the reasoning behind such interactions. Intention, which sums up all interaction information to give the predicted intention of the ego agent, considering its responses to all the interactions. The description of the map environment and agents’ motion status in the dataset are good for training language scene describers, while the interaction and intention parts can be used for training language-based prediction and planning agents.
We claim that to the best of our knowledge, WOMD-Reasoning is the largest real-world language dataset for driving. As shown in Table 1, we provide roughly 3 million Q&A pairs for 63k scenes in WOMD, which is 6-113 times the total Q&As of the datasets we compare to. Furthermore, WOMD-Reasoning contains 346k unique traffic rule-induced and intention-induced interaction Q&As, as well as 63k comprehensive intention predictions of the ego agent containing responses to the interactions, which rarely exist in previous datasets. We will show these features through examples and statistics in the following section.
@@ 4.2 Dataset Quality
WOMD-Reasoning contains traffic rule-induced interaction information Firstly, in Figure 1, we show a traffic scenario of an intersection controlled by the stop sign. In this case, the agent #0 is quite far away from the ego agent at the starting (current) moment, thus their interactions would rarely be considered in previous datasets, as they mainly cover distance-induced interactions. However, the two agents do have a significant interaction: Controlled by the stop signs, the agent #0 has to yield to the ego agent, even though the ego agent is still far from the intersection. This case vividly proves the remarkable importance of the traffic rule-induced interactions we cover.
Another example is a traffic light-controlled intersection shown in Figure 2. Similar to the previous one, agent #0 is initially not close to the ego agent. However, thanks to our analysis of traffic rules, our language dataset effectively captures the interaction that the agent # 1 would yield to the ego agent who is turning left and having right-of-way due to an arrow green light. This further proves that traffic rules are essential in determining the interactions.
Figure 3: A demonstration of Q&As in each part of WOMD-Reasoning and an example for human intention-induced interaction. We show Q&As in all 5 categories regarding the scenario while demonstrating language analysis of overtaking, a human intention-induced interaction in WOMD-Reasoning.
@@
Average ”Accuracy” Rate 80.70% 95.32% 93.51% 94.73% 90.00%
Standard Error 6.08% 2.03% 3.08% 3.51% 10.00%
WOMD-Reasoning contains human intention-induced interaction information Beside traffic rules, our dataset also covers human intention-induced interactions. We provide an additional example to show this, as well as to demonstrate a whole picture of each section of the dataset. In Figure 3, WOMD-Reasoning provides fruitful information on the human intention-induced interaction - overtaking, based on pattern recognition abilities built in the GPT. We also show 1-2 Q&As for each category of the dataset for completeness.
To grab a sense of the accuracy of WOMD-Reasoning, we also execute a preliminary human evaluation. Consent has been obtained from all people interviewed. Note that here we ask evaluators to think a Q&A pair as correct if the correct main answer to the question is included in the answer, so the main answer is reasonably based on the human evaluators’ interpretation of the traffic data. This also does not exclude minor errors (like errors in detailed numbers or very far-away agents), since such errors are hardly distinguishable by the human evaluators. We invite 3 people to judge all Q&As from 11 consecutive scenes of the dataset, which contains 501 Q&As in total. The results are shown in Table 3. We conclude that considering our dataset is fully automatically generated, it achieves a satisfying accuracy, or at least it would significantly reduce the human labor necessary to build a more reliable dataset later. Further human verification of the dataset is still ongoing.
@@ 4.3 Comparison to Existing Datasets
Comparing WOMD-Reasoning with existing datasets, as shown in Table 1, we are the largest real-world dataset, in terms of the total scenes, total Q&As, and in the total interactions included. Also, WOMD-Reasoning supports Q&As in scene descriptions, prediction and planning, making it suitable for training LLMs to perform nearly all major tasks in autonomous driving. Diving even deeper into the interaction reasoning in these datasets, we summarize that most interactions in previous datasets are caused by a very near distance. However, as we show in Sec. 4.2, many interactions happen when the two counterparts are initially far away, but they establish interactions by following the traffic rules or by human intentions. Our WOMD-Reasoning contains a wealth of this interaction information, as shown in Figure 4, the vocabulary statistics of the interaction and intention part of our dataset.
@@
Figure 4: Selected vocabulary statistics of WOMD-Reasoning. We stat vocabularies strongly related to traffic rule-induced and human intention-induced interactions in the interaction and intention part of WOMD-Reasoning, illustrating that our dataset contains an abundance of such interaction descriptions and reasoning.
5 Dataset Effectiveness Validation with Trajectory Prediction Task
In this section, we examine the effectiveness of our language-based interaction analysis with vehicle trajectory prediction tasks. We build a very simple network combining T5 text encoder [20] and Multipath++ [28], seeing significant performance boost when the language is introduced.
Figure 5: Effect of language dataset on trajectory prediction. (a) shows the generated language leads the ego agent to be cautious on the incoming car when turning left, while (b) shows the generated language leads the ego agent not to be too cautious when the other agent is yielding to it.
5.1 Method
Our pipeline is simple, designed for demonstrating the effectiveness of the language dataset. We first follow  [35, 4] to use T5 text encoder [20] to encode the language. Specifically, to show the influence of interaction and intention data, we encode these two parts of language data, followed by a few MLPs to reduce their dimensions. Then, we allow the features of the ego and other agents in Multipath++ pipeline to cross-attend to representations of the ego’s intention language and the other agents’ interactions language respectively. The model is trained end-to-end, with the T5 encoder frozen. The pipeline is illustrated in Fig 6 in Appendix, and implementation details are listed in Sec A.1
Due to the restriction of computational resources, we have not fine-tuned LLMs to generate the interaction language analysis. Therefore, our experiment is only an oracle model method, where the interaction language is directly taken from the dataset with information leaks, i.e. the interaction analysis are imported from the dataset directly. However, this can still prove whether the languages in our dataset are useful, though it remains a question whether it can be generated successfully after fine-tuning the LLM on our dataset.
We perform experiments on both the whole WOMD and part of WOMD with language from WOMD-Reasoning. Note that to ensure consistency, we test the model trained on the whole WOMD training set on the whole WOMD interactive evaluation set, while testing the model trained on the language available part of training set on language available part of evaluation set.
5.2 Quantitative Results
As shown in Table 4, with the introduction of the language, the performance of the trajectory prediction model is substantially enhanced. When training and evaluating on the part of WOMD having language, the language modal helps to reduce the Miss Rate (MR) by $10.14\%$ and the minFDE by $6.90\%$. These significant improvements strongly support the effectiveness of the interaction analysis language in downstream tasks like trajectory predictions.
Table 4: Effectiveness of WOMD-Reasoning on the Prediction Task.
Motion Dataset Metric NOT Using Language Using Language
Whole WOMD $MR_{6}(\downarrow)$ 10.91 10.59
WOMD $\cap$ WOMD-Reasoning $MR_{6}(\downarrow)$ 11.44 10.28
WOMD $\cap$ WOMD-Reasoning $minFDE_{6}(\downarrow)$ 1.16 1.08
5.3 Qualitative Results and Analysis
To thoroughly understand the role that language dataset plays in the language-guided trajectory prediction model, and to verify the correspondence between the language data and the prediction result, we illustrate a few scenarios whose prediction is improved with language. In Figure 5 (a), with the help of language noticing that the ego agent should wait for the traffic light before turning left, the ego agent would be more cautious and its predicted conflict with agent #0 is significantly reduced. Reversely in Figure 5 (b), under the language input saying the ego agent does not need to yield to any other agent, the predicted trajectories do not slow down, thus being closer to the ground truth. Both cases suggest that the prediction results improve consistently with the language.
@@ 6 Conclusion
We build WOMD-Reasoning, a language annotation dataset upon WOMD to support LLMs for analyzing and reasoning interactions. Specifically, we focus on interactions induced by traffic rules and human intentions, such as interactions governed by traffic lights and signs, and interactions like overtaking and following, which are both rarely covered in previous datasets. Besides, our WOMD-Reasoning provides 3 million Q&A pairs, which makes it the largest real-world language dataset in autonomous driving to the best of our knowledge, covering a range of topics from scene descriptions, prediction, to planning. By utilizing our proposed dataset, we significantly improve the performances of the trajectory prediction model Multipath++, where we see matches between the language and the prediction improvements, proving the effectiveness of our interaction analysis for downstream tasks.
Limitations. Although WOMD-Reasoning demonstrates its effectiveness in case studies, human evaluations, and downstream prediction tasks, we only manually verified a small subset. With human verification ongoing, the dataset will achieve even greater reliability. Also, WOMD-Reasoning is subject to the limitations of WOMD. For instance, the raw camera data is not yet publicly available, which restricts the use of WOMD-Reasoning in video-qa. As WOMD continues to release additional modalities such as camera features and lidar data recently, these limitations are expected to be naturally addressed.
Broader Impact. Our work is of broad interest to the natural language processing (NLP) and autonomous driving communities. The proposed approach has no new ethical or social issues on its own, except those inherited from NLP or autonomous driving.
@@ Appendix A Appendix
In this supplementary material, we first show the pipeline we use to introduce language into Multipath++ trajectory prediction model, as well as the training details of Multipath++ on WOMD-Reasoning. Then, we show an example of the interpreted language of the motion data. We also revisit the prompts to build the language dataset. We will make the dataset and codes public.
A.1 Pipeline for Introducing Language into Multipath++
Figure  6 presents the pipeline of introducing language into the Multipath++ trajectory prediction pipeline. We use a batch size of 128 and a learning rate of 1e-4 for Multipath++. For training or evaluating on the whole dataset, some scenarios do not have language, so we input "No captions available for this scene." as the language input. Also even for scenarios with language, language is not available for every agent, so we use "No captions available for this agent." for those agents without language inputs. The t5-v1_1-xxl model is used as the language encoder.
Figure 6: Pipeline of introducing language data into Multipath++
A.2 Example of Translated Motion Scenario
Table  5 presents a part of the interpreted motion scenario by the rule-based program we design.
Descriptions on map environment:
The ego agent is heading towards intersection. The intersection center is 18.0 meters in front of the ego agent, and is 17.0 meters on the left of the ego agent. The intersection is a 4 way intersection. Descriptions on ego agent:
The ego agent is on the 1 lane from the right, out of 3 lanes. Its current speed is 8 m/s. It is accelerating. Traffic Light for the ego agent is red. The ego agent is approaching a crosswalk 3 meters ahead. Descriptions on surrounding (other) agents:
Surrounding agent # 0 is a vehicle. It is 23 meters in front of the ego agent, and is 17 meters on the left of the ego agent. It is heading right of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. It is approaching a crosswalk 3 meters ahead. Surrounding agent # 2 is a vehicle. It is 11 meters on the right of the ego agent, and is 0 meters behind the ego agent. It is heading left of the ego agent. It is not moving. It is on the same side of the intersection as the ego agent. It is 34 meters away from the intersection center.
$\cdots$
Descriptions on agents after 3 seconds:

### v2

@@ 4.1 Dataset Overview
Set Map Ego Agent Other Agents Interactions Intentions Total Scenes
Training 188k 268k 1,635k 287k 52k 2,430k 52k
Validation 41k 58k 341k 59k 11k 510k 11k
@@ 4.2 Dataset Quality
WOMD-Reasoning contains traffic rule-induced interaction information. Firstly, in Figure 1 (a), we show a traffic scenario of an intersection controlled by the stop sign. In this case, the agent #0 is quite far away from the ego agent at the starting (current) moment, thus their interactions would rarely be considered in previous datasets, as they mainly cover distance-induced interactions. However, the two agents do have a significant interaction: Controlled by the stop signs, the agent #0 has to yield to the ego agent, even though the ego agent is still far from the intersection. This case vividly proves the remarkable importance of the traffic rule-induced interactions we cover.
Another example is a traffic light-controlled intersection shown in Figure 1 (b). Similar to the previous one, agent #0 is initially not close to the ego agent. However, thanks to our analysis of traffic rules, our language dataset effectively captures the interaction that the agent #1 would yield to the ego agent who is turning left and having right-of-way due to an arrow green light. This further proves that traffic rules are essential in determining the interactions.
Figure 2: A demonstration of Q&As in each part of WOMD-Reasoning dataset and an example for human intention-induced interaction. We show Q&As in all categories regarding the scenario while demonstrating language analysis of overtaking, a human intention-induced interaction in WOMD-Reasoning dataset.
Figure 3: The Motion-LLaVA pipeline, which fine-tunes a multi-modal model with WOMD-Reasoning. Motion data go through pre-trained motion vector encoders from Multipath++ [51] and a projector layer to serve together with the questions in WOMD-Reasoning as the inputs for the LLM. The answers in WOMD-Reasoning serve as the supervision.
@@
Ave. ‘Accuracy’ 80.70% 95.32% 93.51% 94.73% 90.00%
Std. Error 6.08% 2.03% 3.08% 3.51% 10.00%
WOMD-Reasoning contains human intention-induced interaction information. Beside traffic rules, our dataset also covers human intention-induced interactions. We provide an additional example to show this, as well as to demonstrate a whole picture of each section of the dataset. In Figure 2, WOMD-Reasoning provides fruitful information on the human intention-induced interaction - overtaking, based on pattern recognition abilities built in the GPT. We also show 1-2 Q&As for each category of the dataset for completeness.
To grab a sense of the accuracy of WOMD-Reasoning, we also execute a preliminary human evaluation. Consent was obtained from all the people interviewed. Note that here we ask evaluators to think a Q&A pair as correct if the correct main answer to the question is included in the answer, so the main answer is reasonably based on the human evaluators’ interpretation of the traffic data. This also does not exclude minor errors (like errors in detailed numbers or very far-away agents), since such errors are hardly distinguishable by the human evaluators. We invite 3 people to judge all Q&As from 11 consecutive scenes of the dataset, which contains 501 Q&As in total. The results are shown in Table 3. We conclude that considering our dataset is fully automatically generated, it achieves a satisfying accuracy, or at least it would significantly reduce the human labor necessary to build a more reliable dataset later. Further human verification of the dataset is still ongoing.
4.3 Vision Extension with Simulations
To support applications requiring visual information, we provide an API to extend WOMD-Reasoning scenario descriptions to the vision domain, compensating for the absence of raw camera data in the original WOMD dataset. We achieve this by utilizing an open-source traffic scenario simulation platform, namely ScenarioNet [25], and employing the MetaDrive [24] simulator to replay WOMD-Reasoning scenarios in both BEV layout and ego-view images through 3D rendering. This approach enables visual inputs to support the application of WOMD-Reasoning for future research in the vision and language domain. For video generation, we produce a 90-frame 10 Hz video aligned with Waymo’s trajectory data, including 10 frames of historical footage and 80 frames of future trajectory. Users may choose to use the historical portion for prediction tasks.
@@ 4.4 Comparison to Existing Datasets
Comparing WOMD-Reasoning with existing datasets, as shown in Table 1, we are the largest real-world dataset in terms of the total scenes, total Q&As, and the total interactions included. Also, WOMD-Reasoning supports Q&As in scene descriptions, prediction, and planning. Together with the simulated video of each covered scenario, it is suitable for training LLMs to perform nearly all major tasks in autonomous driving. Diving even deeper into the interaction reasoning in these datasets, we summarize that most interactions in previous datasets are caused by a very near distance. However, as we show in Sec. 4.2, many interactions happen when the two counterparts are initially far away, but they establish interactions by following the traffic rules or by human intentions. Our WOMD-Reasoning contains a wealth of this interaction information, as shown in Figure 4, the vocabulary statistics of the interaction and intention part of our dataset.
5 Experiments with Motion-LLaVA
To demonstrate the benefits of WOMD-Reasoning in applications, in this section, we show the performances of Motion-LLaVA, a multi-modal model fine-tuned on WOMD-Reasoning. We show that the model gains significant abilities from WOMD-Reasoning in answering both factual and reasoning questions related to driving, which is supported by both qualitative showcases of its answers to driving-related questions and quantitative evaluations of language metrics.
5.1 Model Setup
We take LLaVA-v1.5-7b [28] as the pre-trained VLM. Since MultiPath++ [51] did not release their codes nor checkpoints, we pre-train the implementations by [21]. Our multi-modal fine-tuning takes 2 GPU days ($\approx$ 1 day on 2xNVIDIA A6000 GPUs) to train 1 epoch on the entire training set of WOMD-Reasoning. For evaluation, we randomly select $\approx$ 1,000 QA pairs from the validation set of WOMD-Reasoning.
5.2 Qualitative Results
We present a set of Q&A examples answered by Motion-LLaVA covering each category of WOMD-Reasoning in one specific driving scenario, shown in Figure 5. We observe that Motion-LLaVA model is capable of providing accurate narratives and analysis on diverse driving-related contents thanks to the wide coverage of WOMD-Reasoning.
As WOMD-Reasoning is proposed to focus on the interaction analysis, we further present a few interaction analysis Q&As by Motion-LLaVA in Figure 6, in various interaction scenarios. These results further confirm the capacity of Motion-LLaVA for analyzing interactions.
5.3 Benchmarks of Multi-modal Reasoning Models
To comprehensively benchmark our proposed Motion-LLaVA model and fine-tuned LLaVA, we utilize a suite of standard language metrics, including BLEU-4 [35], METEOR [3], CIDEr [52], SPICE [1], and ROUGE-L [27], along with GPT Score [38] to assess the quality of the generated content. Details on the prompts used to calculate the GPT Score can be found in Appendix A.3. We present evaluation metrics of all generated answers, as well as those in the two categories, i.e. the Facts which include the road environment and agents’ motion status, and the Reasonings which cover analysis on interactions and intentions. Benchmark results are shown in Table 4.
@@ 5.3 Benchmarks of Multi-modal Reasoning Models
Figure 4: Selected vocabulary statistics of WOMD-Reasoning. We stat vocabularies strongly related to traffic rule-induced and human intention-induced interactions in the interaction and intention part of WOMD-Reasoning, illustrating that our dataset contains abundant such interaction descriptions and reasoning.
Table 4: Benchmarks of fine-tuned multi-modal models with language metrics.
Questions Model ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) GPT Score ($\uparrow$)
All Shuffled Ground-truths 0.481 0.311 0.263 1.68 0.375 1.48
Non-Fine-tuned LLaVA [28] 0.512 0.211 0.275 1.36 0.455 2.31
Fine-tuned LLaVA [28] 0.779 0.581 0.439 5.51 0.735 6.88
Motion-LLaVA (Ours) 0.792 0.616 0.449 5.69 0.744 7.02
Facts Shuffled Ground-truths 0.528 0.349 0.273 2.26 0.401 2.06
Only Non-Fine-tuned LLaVA [28] 0.561 0.322 0.326 1.50 0.483 2.46
Fine-tuned LLaVA [28] 0.821 0.702 0.495 6.05 0.783 6.91
Motion-LLaVA (Ours) 0.840 0.736 0.516 6.35 0.794 7.09
Reasonings Shuffled Ground-truths 0.479 0.326 0.281 1.81 0.462 1.87
Only Non-Fine-tuned LLaVA [28] 0.262 0.097 0.201 0.06 0.283 0.91
Fine-tuned LLaVA [28] 0.596 0.452 0.357 2.26 0.557 6.14
Motion-LLaVA (Ours) 0.614 0.474 0.366 2.52 0.571 6.76
Figure 5: An example set of answers to diverse driving-related questions generated by Motion-LLaVA on a WOMD scenario. We see that Motion-LLaVA is capable of answering all topics of questions covered by WOMD-Reasoning accurately.
We first verify the effectiveness of these language metrics with shuffled ground-truths, which attach shuffled true answers to questions randomly. Results confirm that low scores are granted for this model, indicating that a similar structure with wrong answer content would not be able to cheat our metrics. Then we benchmark the off-the-shelf LLaVA [28], where BEV plots are used as the vision input for the driving scenario. This non-fine-tuned version earns a very low score in all metrics, confirming that LLaVA itself cannot answer driving-related questions correctly.
Then, we perform fine-tuning on WOMD-Reasoning. We compare two models with different modalities as motion input: LLaVA using the BEV plot of the driving scene as the vision input [28], and Motion-LLaVA which uses motion data as input. We observe that both fine-tuned multimodal models significantly outperform previous baselines across all metrics and question types, underscoring the value of fine-tuning on WOMD-Reasoning for addressing driving-related queries. Compared to the fine-tuned LLaVA, Motion-LLaVA achieves superior performance across all metrics and question types, highlighting the advantages of utilizing the motion data as input. We hypothesize that this performance gap is due to the information loss that occurs when converting motion data into a BEV plot with a scale bar. We also visualize and compare the answers of Motion-LLaVA and these baselines on the same driving scene in Appendix A.4.
Figure 6: An example set of answers to interaction-related questions generated by Motion-LLaVA on various WOMD scenarios. We see that Motion-LLaVA is capable to perform diverse interaction reasonings.
In addition to the language metrics, we provide other objective metrics to evaluate answers to quantitative questions regarding Facts. Of the 1,000 questions in the factual test set, 130 are quantitatively assessable (e.g., position, speed, etc.). For these, we report the root median squared error (Median Error), and accuracy within a 1.0-meter tolerance (Acc1.0), following similar metrics used in [19] as shown in Table 6. Results confirm that Motion-LLaVA is capable of providing accurate answers to factual questions.
Table 5: Benefits of using pre-trained motion vector encoder for fine-tuning multi-modal models on WOMD-Reasoning for factual Q&As.
Motion Vector Encoder Pre-training ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) Acc1.0 ($\uparrow$) Median Error($\downarrow$) GPT Score ($\uparrow$)
None 0.814 0.684 0.491 5.95 0.774 30.0% 2.00 6.79
Motion Caption 0.824 0.711 0.498 6.05 0.775 42.6% 2.24 6.97
AutoEncoder 0.837 0.727 0.516 6.35 0.803 47.3% 1.41 7.07
Motion Prediction (Ours) 0.840 0.736 0.516 6.35 0.794 56.2% 1.00 7.09
Table 6: Accuracy of answers to factual questions.
Model Acc1.0 ($\uparrow$) Median Error ($\downarrow$)
Fine-tuned LLaVA [28] 10.9% 5.0
Motion-LLaVA (Ours) 56.2% 1.0
5.4 Ablation on the Design of Motion-LLaVA
Chain-of-Thought Inference. To address the hallucinated response problem [15, 49, 45] in reasoning interactions, we use a Chain-of-Thought (CoT) approach to provide Motion-LLaVA with structured context for more accurate reasoning. Table 7 compares model performance with and without CoT on interaction and intention analysis. We see that with CoT, the model performs better in GPT Scores. We further show that if our motion data translation program A.1.1 is used during inference to provide contexts, i.e., with factual description context (an example can be found in Appendix A.6) and CoT, the performance can be significantly improved, proving the potential of the CoT inference strategy.
Influence of Pre-training for the Motion Vector Encoder. To establish a pre-training strategy for the motion vector encoder, we examine factual question answering across Motion-LLaVA with four pre-training configurations: no pre-training, autoencoder-style pre-training, motion-caption pre-training [7], and pre-training with a motion prediction task. All configurations shared the same encoder structure. In the motion-caption pre-training, we pre-train all but LLM weights in Motion-LLaVA, by penalizing errors in captioning the motion. For autoencoder-style pre-training and pre-training with a motion prediction task, we retain only the encoder after pre-training, to function as the motion vector encoder for Motion-LLaVA.
The performances of models with these pre-training settings are presented in Table 5. Generally, pre-training motion vector encoders can improve the response quality. Specifically, pre-training with motion prediction tasks outperforms the other two approaches. We attribute this to the close alignment between WOMD-Reasoning ’s question categories and the information needed in motion prediction tasks. As a result, pre-training with the motion prediction task achieves the highest GPT Score and overall accuracy, especially in numerical precision reflected by $Acc_{1.0}$.
@@ 6 Conclusion
We first build WOMD-Reasoning, a language annotation dataset upon WOMD, and then we propose Motion-LLaVA, an LLM fine-tuned on it, both focusing on analyzing and reasoning interactions. Specifically, our WOMD-Reasoning emphasizes interactions induced by traffic rules and human intentions, such as interactions governed by traffic lights and signs, and interactions like overtaking and following, which are both rarely covered in previous datasets. Besides, WOMD-Reasoning provides 3 million Q&A pairs, which makes it the largest real-world language dataset in autonomous driving to the best of our knowledge, covering a range of topics from scene descriptions, prediction, to planning. Boosted by WOMD-Reasoning, we benchmark a set of methods, to propose Motion-LLaVA, which fine-tunes a multi-modal model to accurately answer questions covering both driving scenario descriptions and interaction reasonings. Motion-LLaVA proves the effectiveness of WOMD-Reasoning in assisting real-world driving through various language-involved downstream tasks.
Table 7: Effects of CoT in answering reasoning-related questions.
Reasoning Inference Method ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) GPT Score ($\uparrow$)
Direct 0.616 0.474 0.365 2.54 0.578 6.59
CoT 0.614 0.474 0.366 2.52 0.571 6.76
CoT w/ Facts Contexts 0.664 0.542 0.403 3.10 0.631 7.64
Discussions. Although WOMD-Reasoning demonstrates its effectiveness in case studies, human evaluations, and multi-modal fine-tuning, we only manually verified a small subset. With human verification ongoing, the dataset will achieve even greater reliability. Also, WOMD-Reasoning is subject to the limitations of WOMD, e.g. its accuracy and lack of vision modal. We leverage a simulation approach to address the lack of support for video-Q&A, while acknowledging that it still exhibits a domain gap from real-world visual input. As WOMD continues to release additional modalities such as camera features and lidar data recently, and as the iterative development of the simulator [25, 24] we use, these limitations are expected to be naturally addressed.
@@ Appendix A Appendix
In this appendix, we first show the building process of WOMD-Reasoning in detail in A.1. Specifically, we show the construction of the interpreted language of the motion data with an example (A.1.1), and we revisit the detailed methods utilizing GPT to build the language dataset as well as providing the full set of prompts (A.1.2). Some examples of the vision modal of WOMD-Reasoning provided by simulations are presented in A.2.
Then, for Motion-LLaVA, we show the prompt for obtaining the GPT Score used for evaluation (A.3). We further provide qualitative comparison of benchmarked multi-modal models listed in Figure 5 (A.4). We also provide an example of aggregated context (A.5) as well as factual description context (A.6) used in Chain-of-Thought inference ablations for reference. Finally, an ablation on details in multi-modal model fine-tuning strategy is attached (A.7).
A.1 Details in Building WOMD-Reasoning
In this appendix, we show the details of how we build WOMD-Reasoning dataset, including detailed explanations on each step we take, the example of the translated motion scenarios, and the GPT prompts. The full version of codes and prompts used herein can be found at https://github.com/yhli123/WOMD-Reasoning.
A.1.1 Automatic Translation of Driving Scenarios into Language Descriptions
To ensure enough information for analyzing interactions involving the ego agent, we translate the motion data into an ego-centric description of the traffic scenario. Specifically, before describing the ego agent status, we first narrate the map environment around the ego agent, including information about related intersections, lanes, stop signs, crosswalks, speed bumps, etc. The above information provides a comprehensive spatial guidance that facilitates describing the motion of either ego or other agents accurately. Then, we describe the status of the ego agent in the scenario, including its motion status (e.g., the velocity, acceleration, etc.) and the positional status (e.g., the lane it occupies and its position related to the intersection center). Furthermore, the traffic light and signal information related to it are summarized in the description. Besides the ego agent, the depiction of other agents, including other vehicles, bicycles, and pedestrians, is also crucial for GPT to analyze the interactions. Therefore, we then describe every other agent in the scenario in the same way as we describe the ego agent. We additionally include the positional relation between the other agent and the ego one to support the identification of every agent’s position.
Utilizing the methods above, we can translate motion datasets into languages. Due to the limitations of computational resources of GPT, we only choose to translate those scenarios that have objects_of_interest labels in WOMD, which indicates confirmed significant interactions happening in the scenario. We build two subsets of WOMD-Reasoning with the same setting: the training set is built on the training set of WOMD, while the validation set is built on the interactive validation set of WOMD. In total, we translate 63k scenarios into language. An example of the translation can be found in Table 8.